Jennifer invested $1,000 more in Bond A than she invested in Bond B. Bond A paid 4.15%, while Bond B paid 3.9%. The total yearly interest for both investments was $229.87. How much money did Jennifer invest in Bond A?
Answer by lwsshak3(11628)

Jennifer invested $1,000 more in Bond A than she invested in Bond B. Bond A paid 4.15%, while Bond B paid 3.9%. The total yearly interest for both investments was $229.87. How much money did Jennifer invest in Bond A?
**
let x=money Jennifer invested in Bond B
x+1000=money Jennifer invested in Bond A
..
3.9%x+4.15%(x+1000)=229.87
.039x+.0415x+41.5=229.87
.0805x=188.37
x=188.37/.0805
x=2340
x+1000=3340
money Jennifer invested in Bond A=$3340